# MIRAGE: KV Cache Optimization through Parameter Remapping for Multi-tenant LLM Serving

### v1

@@ Abstract.
KV cache accelerates LLM inference by avoiding redundant computation, at the expense of memory. To support larger KV caches, prior work extends GPU memory with CPU memory via CPU-offloading. This involves swapping KV cache between GPU and CPU memory. However, because the cache updates dynamically, such swapping incurs high CPU memory traffic. We make a key observation that model parameters remain constant during runtime, unlike the dynamically updated KV cache. Building on this, we introduce MIRAGE, which avoids KV cache swapping by remapping, and thereby repurposing, the memory allocated to model parameters for KV cache. This parameter remapping is especially beneficial in multi-tenant environments, where the memory used for the parameters of the inactive models can be more aggressively reclaimed. Exploiting the high CPU–GPU bandwidth offered by the modern hardware, such as the NVIDIA Grace Hopper Superchip, we show that MIRAGE significantly outperforms state-of-the-art solutions, achieving a reduction of 44.8%-82.5% in tail time-between-token latency, 20.7%-99.3% in tail time-to-first-token latency, and 6.6%-86.7% higher throughput compared to vLLM.
copyright: none
@@ 1. Introduction
The memory demands of Large Language Models (LLMs) are growing at a much faster rate than GPU memory capacity, making memory a major bottleneck for efficient inference serving (Hanindhito et al., 2024; Kwon et al., 2023; Touvron et al., 2023b; Dubey et al., 2024; Nvidia, 2025a, d, c). KV cache (Pope et al., 2023; Kwon et al., 2023) was introduced as a significant way to reduce the inference latency: it stores intermediate results from previously generated tokens to avoid recomputing tensor values. To harness the impact of KV cache on the performance of LLM inference serving, numerous memory management techniques are proposed (Pope et al., 2023; Lee et al., 2024a; Liu et al., 2024a; Prabhu et al., 2025; Jin et al., 2024; Adnan et al., 2024), including paging-based mechanisms for runtime KV cache management (Kwon et al., 2023). However, despite these optimizations, when the KV cache size hits the limit of the GPU memory, one must recompute key-value pairs, increasing serving latency (Kwon et al., 2023) (Figure 1). The KV-cache bottleneck increases further in multi-tenant environments where multiple LLMs share and contend over the GPU(s) resources (Yu et al., 2025).
Using CPU memory as an extension of the GPU memory (also called CPU-offloading) is one way to ease the pressure on GPU memory (Peng et al., 2020; Huang et al., 2020; Rajbhandari et al., 2021; Ren et al., 2021; Lim et al., 2021; Sheng et al., 2023; Jiang et al., 2024; Xu et al., 2024). However, it comes at the cost of increased latency or reduced throughput, depending on the CPU-GPU transfer bandwidth. Prior work (Jiang et al., 2024) explored offloading portions of attention computation to CPUs, allowing KV cache to be generated and stored in CPU memory rather than occupying GPU memory. This allowed the CPU–GPU system to handle larger batch sizes; however, limited CPU performance in attention computation can become a bottleneck, potentially increasing token generation latency by about 10-20% in certain scenarios (Jiang et al., 2024).
To make it practical, new advancements from the hardware community have dramatically increased the CPU-GPU data transfer bandwidth: the latest Grace Hopper superchip (GH200) (Nvidia, 2025c) introduces a new NVLink-C2C interface between CPU and GPU, offering 7×7$\times$7 × bandwidth of the H100 (Nvidia, 2025d) and 14×14$\times$14 × that of the A100 (Nvidia, 2025a) (Figure 1). This increased CPU-GPU bandwidth opens new opportunities to design low-latency and high-throughput LLM inference serving systems. At the same time, with the availability of higher bandwidth, the limited parallelism of CPUs, rather than the bandwidth, may emerge as the new performance bottleneck (§ 3.1).
@@
Instead, we propose dynamic parameter remapping [^1] In this paper, swap refers to bidirectional data transfer to and from CPU/GPU memory, whereas remap denotes unidirectional data transfer from CPU to GPU memory., where we simply repurpose a portion of the GPU memory originally allocated for model parameters to support a larger KV cache, enabling the system to accommodate increased KV cache requirements. Unlike KV cache swapping, parameter remapping introduces negligible synchronization overhead, as model parameters remain constant during inference. As shown in Figure 2 (right), it is a non-blocking, unidirectional data transfer that does not require offloading parameters back to the CPU. This immutability simplifies resource management, as the transfer is both deterministic and fixed in size. Notably, parameter remapping is particularly well-suited for multi-tenant LLM serving: it enables efficient and proactive reuse of memory allocated to inactive models [^2] Prior work observed that models frequently enter idle periods without incoming requests. In production traces, over 60% of models exhibit more than 1,000 idle intervals exceeding 10 seconds each (Yu et al., 2025). We refer to models in such intervals as inactive models. by repurposing their parameter storage as KV cache for active models.
@@
However, realizing a practical parameter remapping engine raises several open questions, including when to initiate and stop remapping, which models to target, how many layers to remap, and which specific layers to select. For example, suboptimal layer selection can trigger excessive CPU–GPU data transfers, resulting in a 12.7% reduction in throughput (§ 7.5). Failing to halt remapping in a timely manner can introduce unnecessary pipeline overhead (Hestness et al., 2015; Xu et al., 2024), negatively impacting system performance and increasing serving latency by up to 49% (§ 7.6). The complexity of these decisions is further amplified by the dynamic nature of system load and the variability in request sequence lengths across inputs (Yu et al., 2025; Stojkovic et al., 2024b; Patel et al., 2024). As a result, these decisions must be made adaptively at runtime, highlighting the need for a sophisticated memory management and data transfer engine within the LLM inference serving system (§ 3.4).
@@
In this paper, we introduce MIRAGE, a Dynamic Remapping Engine that automatically and elastically adjusts KV cache size at layer [^3] In this paper, a layer refers to a hidden layer comprising feed-forward, attention, and normalization components, though the exact structure may vary across models (Agrawal et al., 2024a; Liakopoulos et al., 2025; Xu et al., 2024). For example, Llama-2-13B consists of 40 such hidden layers (Touvron et al., 2023b). granularity based on runtime demands. Such flexibility allows MIRAGE to be integrated into any LLM inference serving system with any schedulers (§ 8). We make the following key contributions:
@@
  • We deeply study several potential GPU memory management optimization strategies to motivate dynamic parameter remapping.
@@
  • We introduce MIRAGE, a Dynamic Remapping Engine for multi-tenant LLM inference serving systems that elastically adjusts the KV cache size at runtime by remapping parameter memory as KV cache memory.
@@
  • We integrate MIRAGE into vLLM to evaluate its effectiveness. Compared to the baseline vLLM, MIRAGE increases throughput by 7.9%-86.7% and reduces tail latency by 20.7%-99.3%.
@@ 2.1. LLM Inference Serving
Inference Execution Workflow: LLM inference serving consists of two phases: (a) the prefill phase, which processes the full input sequence in a single pass, and (b) the decode phase, which generates one output token at a time. Figure 3 illustrates the data placement and execution workflow for a typical LLM inference serving system (Kwon et al., 2023). The two primary contributors to GPU memory are LLM model parameters and the KV Cache (Pope et al., 2023; Kwon et al., 2023; Agrawal et al., 2024b; Prabhu et al., 2025). 1 Before the inference serving phase, the model parameters are transferred from CPU memory (DRAM) to GPU memory (HBM) (Li et al., 2023; Kwon et al., 2023; Agrawal et al., 2024b; Prabhu et al., 2025; Fu et al., 2024). After loading parameters, 2 input tokens are then loaded to GPU memory, where 3 tokens are transferred to GPU compute units (where LLM kernel executes) for processing during the prefill phase. During the decode phase, the KV cache is generated and expands with the sequence length. Throughout each iteration, the KV cache is transferred between 4 LLM kernels and 5 GPU memory. When the maximum sequence length is reached or an end-of-sequence (EOS) token is generated, the output tokens produced by the GPU compute units are first written to GPU memory 6, and then transferred back to CPU memory 7.
@@
Performance Metrics: The performance of an LLM inference serving system is usually evaluated using two key latency metrics: TTFT (time-to-first-token) and TBT (time-between-tokens). TTFT measures the time taken to generate the first output token from the moment a request enters the system, reflecting the initial responsiveness of an inference serving system. TBT measures the interval between generating consecutive output tokens for a given request, which reflects the overall perceived smoothness of the response. Throughput is another critical metric to evaluate inference serving systems alongside latency, which quantifies the number of requests/tokens a system can handle within a specific time frame.
@@
2.2. Memory Bottlenecks in LLM Serving
The memory demands of LLMs have increased at a much faster rate than GPU memory capacity, making memory a critical bottleneck in LLM inference serving. As model sizes grow, they require not only greater computational resources but also significantly more memory to accommodate both model parameters and KV cache (Kwon et al., 2023; Agrawal et al., 2024b). To address this challenge, recent works focused on efficiently managing KV cache, as its size grows linearly with the sequence length and batch size, which are unknown prior to runtime (Pope et al., 2023; Kwon et al., 2023; Lee et al., 2024a; Liu et al., 2024a; Prabhu et al., 2025; Jin et al., 2024; Adnan et al., 2024). For example, vLLM employs PagedAttention (Kwon et al., 2023), which leverages a paging mechanism to manage KV cache pages during runtime. However, when the KV cache exceeds the available GPU memory, recomputation becomes necessary. This scenario is difficult to anticipate, as the sequence length is unknown before runtime and request arrival patterns vary in production environments (Stojkovic et al., 2024a; Patel et al., 2024; Stojkovic et al., 2024b). This unpredictability also affects multi-tenant LLM inference systems. Regardless of the scheduling policy, systems using PagedAttention face the same challenge as single-model serving: runtime KV cache usage is inherently unpredictable. Recomputation can significantly degrade the performance of the affected sequence, especially by increasing the tail latency of the overall serving system. To mitigate this, a mechanism is desired to prevent recomputation when KV cache capacity is exceeded, thereby reducing tail latency in LLM inference serving.
@@ 3. Motivation
This section presents three approaches to expanding KV cache capacity, enabled by the significantly higher CPU–GPU bandwidth of modern hardware architectures like GH200.
  • Approach 1: offloading computation to CPUs (Jiang et al., 2024), where a portion of the attention operations is offloaded to CPUs, allowing the KV cache to be generated and stored in CPU memory instead of consuming GPU memory. As a result, the combined CPU–GPU system can accommodate larger batch sizes, leading to increased throughput (§ 3.1).
@@ 3.1. Limitations of CPU Offloading
Prior works have demonstrated that offloading part of the attention computation and KV cache states from GPUs to CPUs can effectively increase the batch size and thus inference throughput in LLM inference serving systems (Jiang et al., 2024). Although offloading computation to CPUs has shown promising performance on earlier hardware platforms, we find it is less efficient on advanced systems such as the GH200 superchip, where limited CPU parallelism, rather than CPU-GPU bandwidth, is the bottleneck. Figure 5 presents the CPU computation time for processing OPT-13b across varying batch sizes serving requests from the ShareGPT dataset (Team, 2025). We compare two approaches: (a) offloading computations to CPUs, and (b) continuing GPU computation while remapping parameters to expand KV cache capacity to support large batch sizes. Parameter remapping is preferable when the parameter loading time does not exceed the CPU computation time. In GH200 systems, high CPU–GPU bandwidth but limited CPU parallelism make parameter loading significantly faster than CPU processing, reinforcing the suitability of parameter remapping. In contrast, on an H100, offloading computation to CPUs may be viable, even when compared to a low remapping percentage [^4] We define the remapping percentage as the fraction of memory used by model parameters that are remapped as added capacity for KV cache, while the remaining portion continues to be used for storing model parameters., such as 30%. Therefore, on systems with high CPU-GPU bandwidth, remapping parameters offers a more promising solution than offloading computation to CPUs.

### v2

# Oneiros: KV Cache Optimization through Parameter Remapping for Multi-tenant LLM Serving

@@ Abstract.
KV cache accelerates LLM inference by avoiding redundant computation, at the expense of memory. To support larger KV caches, prior work extends GPU memory with CPU memory via CPU-offloading. This involves swapping KV cache between GPU and CPU memory. However, because the cache updates dynamically, such swapping incurs high CPU memory traffic. We make a key observation that model parameters remain constant during runtime, unlike the dynamically updated KV cache. Building on this, we introduce Oneiros, which avoids KV cache swapping by remapping, and thereby repurposing, the memory allocated to model parameters for KV cache. This parameter remapping is especially beneficial in multi-tenant environments, where the memory used for the parameters of the inactive models can be more aggressively reclaimed. Exploiting the high CPU-GPU bandwidth offered by the modern hardware, such as the NVIDIA Grace Hopper Superchip, we show that Oneiros significantly outperforms state-of-the-art solutions, achieving a reduction of 44.8%-82.5% in tail time-between-token latency, 20.7%-99.3% in tail time-to-first-token latency, and 6.6%-86.7% higher throughput compared to vLLM. Source code of Oneiros is available at https://github.com/UT-SysML/Oneiros/.
Large language models, Multi-tenant inference serving systems
journalyear: 2025copyright: ccconference: ACM Symposium on Cloud Computing; November 19–21, 2025; Online, USAbooktitle: ACM Symposium on Cloud Computing (SoCC ’25), November 19–21, 2025, Online, USAdoi: 10.1145/3772052.3772215isbn: 979-8-4007-2276-9/2025/11ccs: Computer systems organization Cloud computingccs: Computing methodologies Machine learningccs: Software and its engineering Memory management
@@ 1. Introduction
Equal contribution.
The memory demands of Large Language Models (LLMs) are growing at a much faster rate than GPU memory capacity, making memory a major bottleneck for efficient inference serving (Hanindhito et al., 2024; Kwon et al., 2023; Touvron et al., 2023b; Dubey et al., 2024; Nvidia, 2025a, d, c). KV cache (Pope et al., 2023; Kwon et al., 2023) was introduced as a key mechanism to reduce inference latency: it stores intermediate results from previously generated tokens to avoid recomputing tensor values. To harness the impact of KV cache on the performance of LLM inference serving, numerous memory management techniques are proposed (Pope et al., 2023; Lee et al., 2024a; Liu et al., 2024; Prabhu et al., 2025; Jin et al., 2024; Adnan et al., 2024), including paging-based mechanisms for runtime KV cache management (Kwon et al., 2023). However, despite these optimizations, when the KV cache size hits the limit of the GPU memory, one must recompute key-value pairs, increasing serving latency (Kwon et al., 2023) (Figure 1). The KV-cache bottleneck increases further in multi-tenant environments where multiple LLMs share and contend over the GPU(s) resources (Yu et al., 2025).
Using CPU memory as an extension of the GPU memory (also called CPU-offloading) is one way to ease the pressure on GPU memory (Peng et al., 2020; Huang et al., 2020; Rajbhandari et al., 2021; Ren et al., 2021; Lim et al., 2021; Sheng et al., 2023; Jiang et al., 2025; Xu et al., 2024; Na et al., 2025). However, it comes at the cost of increased latency or reduced throughput, depending on the CPU-GPU transfer bandwidth. Prior work (Jiang et al., 2025) explored offloading portions of attention computation to CPUs, allowing KV cache to be generated and stored in CPU memory rather than occupying GPU memory. This allowed the CPU-GPU system to handle larger batch sizes; however, limited CPU performance in attention computation can become a bottleneck, potentially increasing token generation latency by about 10-20% in certain scenarios (Jiang et al., 2025).
To make it practical, new advancements from the hardware community have dramatically increased the CPU-GPU data transfer bandwidth: the latest Grace Hopper superchip (GH200) (Nvidia, 2025c) introduces a new NVLink-C2C interface between CPU and GPU, offering $7\times$ bandwidth of the H100 (Nvidia, 2025d) and $14\times$ that of the A100 (Nvidia, 2025a) (Figure 1). This increased CPU-GPU bandwidth opens new opportunities to design low-latency and high-throughput LLM inference serving systems. At the same time, with the availability of higher bandwidth, the limited parallelism of CPUs, rather than the bandwidth, may emerge as the new performance bottleneck (§ 3.1).
@@
Instead, we propose dynamic parameter remapping [^1] In this paper, swap refers to bidirectional data transfer to and from CPU/GPU memory, whereas remap denotes unidirectional data transfer from CPU to GPU memory., where we simply repurpose a portion of the GPU memory originally allocated for model parameters to support a larger KV cache, enabling the system to accommodate increased KV cache requirements. Unlike KV cache swapping, parameter remapping introduces negligible synchronization overhead, as model parameters remain constant during inference. As shown in Figure 2 (right), it is a non-blocking, unidirectional data transfer that does not require offloading parameters back to the CPU. This immutability simplifies resource management, as the transfer is both deterministic and fixed in size. Notably, parameter remapping is particularly well-suited for multi-tenant LLM serving: it enables efficient and proactive reuse of memory allocated to inactive models [^2] Prior work observed that models frequently enter idle periods without incoming requests (Yu et al., 2025). We refer to models in such intervals as inactive models. by repurposing their parameter storage as KV cache for active models.
@@
However, realizing a practical parameter remapping engine raises several open questions, including when to initiate and stop remapping, which models to target, how many layers to remap, and which specific layers to select. For example, suboptimal layer selection can trigger excessive CPU-GPU data transfers, resulting in a 12.7% reduction in throughput (§ 7.5). Failing to halt remapping in a timely manner can introduce unnecessary pipeline overhead (Hestness et al., 2015; Xu et al., 2024), negatively impacting system performance and increasing serving latency by up to 49% (§ 7.6). The complexity of these decisions is further amplified by the dynamic nature of system load and the variability in request sequence lengths across inputs (Yu et al., 2025; Stojkovic et al., 2024b; Patel et al., 2024). As a result, these decisions must be made adaptively at runtime, highlighting the need for a sophisticated memory management and data transfer engine within the LLM inference serving system (§ 3.4).
@@
In this paper, we introduce Oneiros, a Dynamic Remapping Engine that automatically and elastically adjusts KV cache size at layer granularity based on runtime demands. Such flexibility allows Oneiros to be integrated into any LLM inference serving system with any schedulers (§ 8). We make the following key contributions:
@@
  • We study several GPU memory management optimization strategies to motivate dynamic parameter remapping.
@@
  • We introduce Oneiros, a Dynamic Remapping Engine for multi-tenant LLM inference serving systems that elastically adjusts the KV cache size at runtime by remapping parameter memory as KV cache memory.
@@
  • We integrate Oneiros into vLLM to evaluate its effectiveness. Compared to the baseline vLLM, Oneiros increases throughput by 7.9%-86.7% and reduces tail latency by 20.7%-99.3%.
@@ 2.1. LLM Inference Serving
Inference Execution Workflow: LLM inference serving consists of two phases: (a) the prefill phase, which processes the full input sequence in a single pass, and (b) the decode phase, which generates one output token at a time. Figure 3 illustrates the data placement and execution workflow for a typical LLM inference serving system (Kwon et al., 2023). The two primary contributors to GPU memory are LLM model parameters and the KV Cache (Pope et al., 2023; Kwon et al., 2023; Agrawal et al., 2024; Prabhu et al., 2025). 1 Before the inference serving phase, the model parameters are transferred from CPU memory (DRAM) to GPU memory (HBM) (Li et al., 2023; Kwon et al., 2023; Agrawal et al., 2024; Prabhu et al., 2025; Fu et al., 2024). After loading parameters, 2 input tokens are then loaded into GPU memory, where 3 tokens are transferred to GPU compute units (where LLM kernel executes) for processing during the prefill phase. During the decode phase, the KV cache is generated and expands with the sequence length. Throughout each iteration, the KV cache is transferred between 4 LLM kernels and 5 GPU memory. When the maximum sequence length is reached or an end-of-sequence (EOS) token is generated, the output tokens produced by the GPU compute units are first written to GPU memory 6, and then transferred back to CPU memory 7.
@@
Performance Metrics: The performance of an LLM inference serving system is typically characterized by two latency metrics –TTFT (time-to-first-token) and TBT (time-between-tokens) – and by throughput. TTFT measures the delay before the first token is generated, indicating system responsiveness, while TBT captures the interval between consecutive tokens, reflecting response smoothness. Throughput quantifies how many requests or tokens the system can process per unit time.
2.2. Memory Bottlenecks in LLM Serving
The memory demands of LLMs have increased at a much faster rate than GPU memory capacity, making memory a critical bottleneck in LLM inference serving. As model sizes grow, they require not only greater computational resources but also significantly more memory to accommodate both model parameters and KV cache (Kwon et al., 2023; Agrawal et al., 2024). To address this challenge, recent works focused on efficiently managing KV cache, as its size grows linearly with the sequence length and batch size, which are unknown prior to runtime (Pope et al., 2023; Kwon et al., 2023; Lee et al., 2024a; Liu et al., 2024; Prabhu et al., 2025; Jin et al., 2024; Adnan et al., 2024). For example, vLLM employs PagedAttention (Kwon et al., 2023), which leverages a paging mechanism to manage KV cache pages during runtime. However, when the KV cache exceeds the available GPU memory, recomputation becomes necessary. This scenario is difficult to anticipate, as the sequence length is unknown before runtime and request arrival patterns vary in production environments (Stojkovic et al., 2024a; Patel et al., 2024; Stojkovic et al., 2024b). This unpredictability also affects multi-tenant LLM inference systems. Regardless of the scheduling policy, systems using PagedAttention face the same challenge as single-model serving: runtime KV cache usage is inherently unpredictable. Recomputation can significantly degrade the performance of the affected sequence, especially by increasing the tail latency of the overall serving system. To mitigate this, a mechanism is desired to prevent recomputation when KV cache capacity is exceeded, thereby reducing tail latency in LLM inference serving.
@@ 3. Motivation
This section presents three approaches to expanding KV cache capacity, enabled by the significantly higher CPU-GPU bandwidth of modern hardware architectures like GH200.
  • Approach 1: offloading computation to CPUs (Jiang et al., 2025), where a portion of the attention operations is offloaded to CPUs, allowing the KV cache to be generated and stored in CPU memory instead of consuming GPU memory. As a result, the combined CPU-GPU system can accommodate larger batch sizes, leading to increased throughput (§ 3.1).
@@ 3.1. Limitations of CPU Offloading
Prior works have demonstrated that offloading part of the attention computation and KV cache states from GPUs to CPUs can effectively increase the batch size and thus inference throughput in LLM inference serving systems (Jiang et al., 2025). Although offloading computation to CPUs has shown promising performance on earlier hardware platforms, we find it is less efficient on advanced systems such as the GH200 superchip, where limited CPU parallelism, rather than CPU-GPU bandwidth, is the bottleneck. Figure 4 presents the CPU computation time for processing OPT-13b across varying batch sizes serving requests from the ShareGPT dataset (Team, 2025). We compare two approaches: (a) offloading computations to CPUs, and (b) continuing GPU computation while remapping parameters to expand KV cache capacity to support large batch sizes. Parameter remapping is preferable when the parameter loading time does not exceed the CPU computation time. In GH200 systems, high CPU-GPU bandwidth but limited CPU parallelism make parameter loading significantly faster than CPU processing, reinforcing the suitability of parameter remapping. In contrast, on an H100 system, offloading computation to CPUs may be viable, even when compared to a low remapping percentage (the fraction of memory used by model parameters that are remapped as added capacity for KV cache), such as 30%. Therefore, on systems with high CPU-GPU bandwidth, remapping parameters offers a more promising solution than offloading computation to CPUs.
(a) CPU Compute.
(b) GPU Compute.
3.2. Limitations of Swapping KV Cache
With emerging hardware offering higher CPU-GPU bandwidth, KV cache for non-executing (future) layers can be temporarily swapped into CPU memory and brought back to GPU memory when needed. This approach enables larger batch sizes and enhances overall system performance, as demonstrated by Pie (Xu et al., 2024). However, KV cache swapping presents significant challenges, primarily because the KV cache must be updated after every token generation (Yu et al., 2022; Kwon et al., 2023; Agrawal et al., 2024). This results in frequent synchronization between CPU and GPU memory, introducing overhead that affects serving latency. In addition to synchronization overhead, the bidirectional nature of data transfers can reduce effective CPU-GPU bandwidth, preventing the system from achieving an ideal pipeline of data transfer and computation. Our measurements of CPU host memory (LPDDR5X [^3] The trend observed in our measurements also holds for other memory technologies (e.g., DDR4, DDR5) beyond LPDDR5X (Esmaili-Dokht et al., 2024).) bandwidth show that unidirectional (read-only of CPU host memory) CPU-to-GPU transfers can reach $\sim$427 GB/s. However, when the read-to-write ratio is 1:1, bandwidth drops to $\sim$366 GB/s, resulting in a $\sim$15% reduction. This bandwidth degradation can further impact serving throughput.
3.3. Our Proposal: Parameter Remapping
We introduce Parameter Remapping, where, instead of swapping KV cache, we propose to dynamically remap the space used for model parameters to automatically and elastically expand the KV cache size at runtime. Dynamic parameter remapping reclaims GPU memory by repurposing a portion of the model parameters for KV cache usage during runtime, enabling online inference serving systems to handle peak loads when GPU memory becomes insufficient for KV cache. As shown in the two examples in Figure 5, deploying two OPT-13b models on an H100 GPU using either temporal (Figure 5) or spatial (Figure 5) sharing one GPU consumes approximately 52GB of the available 80GB memory for model parameters (Kwon et al., 2023), leaving only about 28GB for KV cache. By remapping one-third of the model parameters as KV Cache – either $2/3$ from the inactive model in temporal sharing or $1/3$ from each model in spatial sharing – the maximum batch size can be increased by up to $1.75\times$, resulting in a corresponding $1.75\times$ improvement in system throughput. Unlike KV cache swapping, parameter remapping involves non-blocking, unidirectional data transfers and does not require backing up model parameters from GPU to CPU, as parameters remain unchanged during inference. As a result, synchronization overhead is significantly reduced. Additionally, parameter remapping is particularly effective in multi-tenant LLM inference scenarios, such as multi-agent workflows (Talebirad and Nadiri, 2023; Chan et al., 2023; Li et al., 2024b; Han et al., 2024b; Kim et al., 2025) and production deployments where many models remain idle for long periods of time (Yu et al., 2025). In such settings, memory allocated to inactive models can be reclaimed for active ones, a flexibility that KV cache swapping lacks, as inactive models may have no KV cache to be swapped.